### v3

# Joint Discriminative-Generative Modeling via Dual Adversarial Training

Abstract
Simultaneously achieving robust classification and high-fidelity generative modeling within a single framework presents a significant challenge. Hybrid approaches, such as Joint Energy-Based Models (JEM), interpret classifiers as EBMs but are often limited by the instability and poor sample quality inherent in Stochastic Gradient Langevin Dynamics (SGLD)-based training. We address these limitations by proposing a novel training framework that integrates adversarial training (AT) principles for both discriminative robustness and stable generative learning. The proposed method introduces three key innovations: (1) the replacement of SGLD-based JEM learning with a stable, AT-based approach that optimizes the energy function by discriminating between real data and Projected Gradient Descent (PGD)-generated contrastive samples using the BCE loss; (2) synergistic adversarial training for the discriminative component that enhances classification robustness while eliminating the need for explicit gradient penalties; and (3) a two-stage training strategy that addresses normalization-related instabilities and enables leveraging pretrained robust classifiers, generalizing effectively across diverse architectures. Experiments on CIFAR-10/100 and ImageNet demonstrate that our approach: (1) is the first EBM-based hybrid to scale to high-resolution datasets with high training stability, simultaneously achieving state-of-the-art discriminative and generative performance on ImageNet $\times256$; (2) uniquely combines generative quality with adversarial robustness, enabling critical applications like robust counterfactual explanations; and (3) functions as a competitive standalone generative model, matching the generative quality of autoregressive methods (VAR-d16) and surpassing diffusion models while offering unique versatility.
1 Introduction
Deep learning models have traditionally been developed with either discriminative or generative objectives in mind, rarely excelling at both simultaneously (Ng and Jordan, 2001; Jebara, 2004; Lasserre et al., 2006; Xie et al., 2016; Grathwohl et al., 2019). Discriminative models are optimized for classification or regression tasks but lack the ability to model data distributions, while generative models can synthesize new data samples but may underperform on downstream classification tasks (Ng and Jordan, 2001; Jebara, 2004). Recent research has explored unifying these approaches through joint discriminative-generative modeling frameworks that aim to combine the predictive power of discriminative approaches with the rich data understanding of generative models (Xie et al., 2016; Lazarow et al., 2017; Jin et al., 2017; Du and Mordatch, 2019; Grathwohl et al., 2019; Chen et al., 2019; Guo et al., 2023; Deja et al., 2023).
Among these unification efforts, Energy-Based Models (EBMs) have emerged as a promising framework due to their flexibility and theoretical connections to both paradigms. In particular, Joint Energy-Based Models (JEM) (Grathwohl et al., 2019) demonstrated that standard classifier architectures could be reinterpreted to simultaneously function as EBMs, enabling both high-accuracy classification and reasonable sample generation. However, a critical limitation of JEM and similar approaches is their reliance on Markov Chain Monte Carlo (MCMC) methods such as Stochastic Gradient Langevin Dynamics (SGLD) for training the generative component. SGLD-based EBM learning suffers from significant training instabilities, computational inefficiency, and often produces poor-quality samples (Grathwohl et al., 2019; Duvenaud et al., 2021; Du and Mordatch, 2019; Zhao et al., 2020; Gao et al., 2018; Nijkamp et al., 2019), limiting the practical adoption of these hybrid models.
We address these limitations by introducing Dual Adversarial Training (DAT), a novel framework that leverages adversarial training (AT) principles for both discriminative robustness and stable generative learning within a unified JEM-based architecture. Our approach employs a dual application of adversarial training: (1) standard AT for the discriminative component to achieve robustness against adversarial perturbations, and (2) an AT-based energy function learning strategy for the generative component that replaces unstable SGLD-based JEM learning.
Our key technical contributions include:
  1. 1.
    A stable AT-based alternative to SGLD-based JEM learning. We replace the unstable SGLD-based JEM learning with an adversarial training approach that optimizes the energy function through Binary Cross-Entropy loss using Projected Gradient Descent (PGD; Madry et al. 2017)-generated contrastive samples. This fundamentally addresses the training instabilities that have plagued JEM, enabling reliable convergence and significantly improved sample quality.
  2. 2.
    Adversarial training with synergistic effects. We incorporate adversarial training for the discriminative component, which not only enhances classification robustness but also eliminates the need for explicit $R_{1}$ gradient penalty required by previous AT-EBMs frameworks (Yin et al., 2022), simplifying the training procedure and avoiding constraints on model expressiveness.
  3. 3.
    Two-stage training strategy. We introduce a two-stage training strategy that provides practical benefits including leveraging pretrained robust classifiers and addressing normalization-related instabilities. This strategy generalizes across diverse architectures, working effectively for both ResNet (with batch normalization) and ConvNeXt (with layer normalization).
Experiments on CIFAR-10/100 and ImageNet demonstrate the effectiveness and scalability of our approach. Notably, our approach represents the first hybrid model to achieve competitive generative quality, strong classification performance, and adversarial robustness simultaneously—establishing that hybrid models need not compromise on any dimension. This establishes three advances for hybrid modeling:
  1. 1.
    First EBM-based hybrid that scales to high-resolution complex datasets. Prior EBM-based hybrid approaches (IGEBM, JEM, SADA-JEM) could not scale beyond low resolution or achieve competitive generative performance on ImageNet-level datasets due to SGLD instability. Our approach is the first to overcome these limitations, achieving competitive generative quality and strong classification performance on ImageNet $\times256$ with high training stability, demonstrating that EBM-based hybrid models can scale reliably to complex, high-resolution datasets.
  2. 2.
    Generative quality and adversarial robustness for critical applications. While other scalable hybrids like the diffusion-based EGC achieve reasonable performance, they do not reach state-of-the-art generative quality and lack adversarial robustness. Our dual capability of generative quality and adversarial robustness enables robust counterfactual explanations that are substantially more faithful to target class characteristics than non-robust or robustness-only methods (Section˜4.3.2), essential for interpretable ML in high-stakes domains.
  3. 3.
    Flexibility as a competitive generative model. When evaluated purely on generation quality, our approach (with ConvNeXt-Large) matches the autoregressive model VAR-d16 and surpasses diffusion models (ADM, LDM) on ImageNet $\times256$, while achieving significantly higher throughput than diffusion models. Beyond this competitive quality, our energy-based approach offers unique versatility for diverse synthesis tasks (inpainting, super-resolution, image manipulation) and compositional generation (Du et al., 2020; Santurkar et al., 2019).
These results demonstrate that hybrid models are no longer a compromise in either generative quality or classification performance—they can match state-of-the-art specialized models in both dimensions while providing the unique capabilities described above.
2 Related work
Joint discriminative-generative modeling The pursuit of joint discriminative-generative modeling, or hybrid modeling, aims to combine the predictive power of discriminative approaches with the rich data understanding of generative models within a single framework. This line of research is motivated by the potential to improve classifier robustness, calibration, and out-of-distribution detection (Grathwohl et al., 2019; Du and Mordatch, 2019), while also enabling tasks like sample generation (e.g., for counterfactual explanation (Deja et al., 2023)) and semi-supervised learning (Kingma et al., 2014). A significant thrust in this area involves Energy-Based Models (EBMs). Early work by Xie et al. (2016) showed how generative ConvNets could be derived from discriminative ones, framing them as EBMs. Du and Mordatch (2019) demonstrated that implicitly generative EBMs can achieve strong performance on discriminative tasks like adversarially robust classification and out-of-distribution detection, while addressing scalable EBM training challenges. Grathwohl et al. (2019) introduced Joint Energy-Based Models (JEM), which explicitly reinterpret standard classifiers as EBMs over the joint distribution of data and labels $p(x,y)$, allowing simultaneous classification and generation. Yang et al. (2023) incorporated sharpness-aware minimization (SAM) to smooth energy landscapes and removed data augmentation from the EBM loss term to improve both classification accuracy and generation quality of JEM. Guo et al. (2023) proposed EGC, which employs Fisher divergence within a diffusion framework to learn an unconditional score function $\nabla\log p(x)$ and a conditional classifier $p(y|x)$ for unified classification and generation, thereby circumventing the computational challenges of traditional energy-based model training.
Alternative architectural approaches have also been explored for joint modeling. Rather than energy-based formulations, joint diffusion models (Deja et al., 2023) attach classifiers directly to diffusion model UNet encoders for joint end-to-end training. Another distinct approach is “introspective learning,” where a single model functions as both a generator and a discriminator through an iterative self-evaluation process, developed across works by Lazarow et al. (2017), Jin et al. (2017), and Lee et al. (2018). Flow-based models have also been explored for hybrid tasks; for instance, Residual Flows (Chen et al., 2019) utilized invertible ResNet and showed competitive performance in joint generative and discriminative settings, offering an alternative to EBMs by allowing exact likelihood computation. These diverse approaches underscore the continued effort to create models that synergistically leverage both discriminative and generative learning.
Joint Energy-Based Models (JEM) A significant step towards unifying discriminative and generative modeling within a single framework was presented by Grathwohl et al. (2019) with their Joint Energy-Based Model (JEM). Their key insight was to reinterpret the logits produced by a standard discriminative classifier, typically used to model $p(y|x)$, as defining an energy function for the joint distribution $p(x,y)$. Specifically, they defined the energy $E_{\theta}(x,y)$ as the negative of the logit corresponding to class $y$, $E_{\theta}(x,y)=-f_{\theta}(x)[y]$. This formulation allows for the recovery of the standard conditional distribution $p(y|x)$ via softmax normalization over $y$, while also yielding an unnormalized probability density $p(x)$ by marginalizing out $y$, effectively using the negative LogSumExp of the logits as the energy function for $p(x)$. They proposed a hybrid training objective that combines the standard cross-entropy loss for $p(y|x)$ with an EBM-based objective for $p(x)$ optimized using Stochastic Gradient Langevin Dynamics (SGLD) (Welling and Teh, 2011). Grathwohl et al. (2019) demonstrated that this joint training approach allows JEM to achieve strong performance on both classification and generative tasks, while simultaneously improving classifier calibration, out-of-distribution detection capabilities, and robustness against adversarial examples compared to standard discriminative training.
Our work builds upon these foundations by incorporating adversarial training principles into the joint modeling framework. Our approach is motivated by the theoretical insight that adversarially robust classifiers implicitly learn energy functions (Zhu et al., 2021; Wang et al., 2022b; Mirza et al., 2024), and we draw particularly from recent advances in adversarial training for EBMs (Yin et al., 2022) and methods for achieving robustness on both in-distribution and out-of-distribution data (Augustin et al., 2020; Korst and Asadulaev, 2022). Applications of robust classifiers to image synthesis are explored by Santurkar et al. (2019), who demonstrate that robust classifiers can perform various image synthesis tasks through gradient-based optimization. For a comprehensive discussion of these approaches, see Appendix A.1.
3 Method
3.1 Joint Energy-Based Model
Our approach builds upon the Joint Energy-Based Model (JEM) framework introduced by Grathwohl et al. (2019), which reinterprets the outputs of a standard discriminative classifier as an energy-based model (EBM) over the joint distribution of data $x$ and labels $y$. Given a classifier network that produces logits $f_{\theta}(x)\in\mathbb{R}^{K}$ for $K$ classes, JEM defines the joint energy function as:
$E_{\theta}(x,y)=-f_{\theta}(x)[y]$ (1)
where $f_{\theta}(x)[y]$ is the logit corresponding to class $y$. This energy function can be normalized to obtain a joint probability density:
$p_{\theta}(x,y)=\frac{\exp(-E_{\theta}(x,y))}{Z(\theta)}=\frac{\exp(f_{\theta}(x)[y])}{Z(\theta)}$ (2)
where $Z(\theta)=\sum_{y^{\prime}}\int\exp(f_{\theta}(x^{\prime})[y^{\prime}])dx^{\prime}$ is the partition function (an intractable global normalizing constant). By marginalizing out the label $y$, a marginal density over the input data $x$ can be obtained:
$p_{\theta}(x)=\sum_{y}p_{\theta}(x,y)=\frac{\sum_{y}\exp(f_{\theta}(x)[y])}{Z(\theta)}$ (3)
Thus, a valid energy function for $p_{\theta}(x)$ is given by:
$E_{\theta}(x)=-\log\sum_{y}\exp(f_{\theta}(x)[y])$ (4)
This energy is related to the marginal density by $p_{\theta}(x)=\frac{\exp(-E_{\theta}(x))}{Z(\theta)}$.
A JEM is trained by maximizing the joint log-likelihood $\log p_{\theta}(x,y)$ over labeled training datapoints $(x,y)$ drawn from an empirical joint distribution $p_{\text{data}}(x,y)$. The joint log-likelihood is typically factorized as $\log p_{\theta}(y|x)+\log p_{\theta}(x)$. The conditional term $\log p_{\theta}(y|x)$ can be maximized by minimizing the standard cross-entropy classification loss. The marginal term $\log p_{\theta}(x)$ is optimized using the EBM gradient (LeCun et al., 2006):
$\nabla_{\theta}\mathbb{E}_{x\sim p_{\text{data}}(x)}[\log p_{\theta}(x)]=\mathbb{E}_{x\sim p_{\text{data}}(x)}[-\nabla_{\theta}E_{\theta}(x)]-\mathbb{E}_{x\sim p_{\theta}(x)}[-\nabla_{\theta}E_{\theta}(x)]$ (5)
where $p_{\text{data}}(x)$ is the empirical marginal distribution obtained by marginalizing $y$ from $p_{\text{data}}(x,y)$. This gradient decreases the energy of real data samples while increasing the energy of model-generated samples. At equilibrium when $p_{\theta}(x)=p_{\text{data}}(x)$, these terms balance and the gradient becomes zero.
To approximate the expectation $\mathbb{E}_{x\sim p_{\theta}(x)}[\cdot]$, samples are drawn from $p_{\theta}(x)$ using Stochastic Gradient Langevin Dynamics (SGLD) (Welling and Teh, 2011). SGLD generates samples $x$ starting from an initial distribution $p_{0}(x)$ (e.g., uniform noise) and iteratively applies the update rule:
$x_{t+1}=x_{t}-\frac{\alpha}{2}\nabla_{x}E_{\theta}(x_{t})+\xi_{t},\quad\text{where }\xi_{t}\sim\mathcal{N}(0,\alpha)$ (6)
Here, $\alpha$ is the step size, and $\nabla_{x}E_{\theta}(x_{t})$ is the gradient with respect to the marginal energy function.
3.2 Learning JEM with adversarial training
The JEM framework successfully integrates generative modeling into classifiers, but its reliance on SGLD and EBM gradient (eq.˜5) causes significant training instabilities (Grathwohl et al., 2019; Duvenaud et al., 2021) and results in poor sample quality. We address these limitations by replacing the SGLD-based JEM with an adversarial training (AT) approach inspired by AT-EBMs (Yin et al., 2022).
Specifically, we replace the standard EBM gradient (eq.˜5) with a stabilized formulation:
$\mathbb{E}_{x\sim p_{\text{data}}(x)}[-\nabla_{\theta}E_{\theta}(x)]-\mathbb{E}_{x\sim p_{\theta}(x)}[-\nabla_{\theta}E_{\theta}(x)]$
$\Longrightarrow\;\mathbb{E}_{x\sim p_{\text{data}}(x)}[-\alpha(x)\nabla_{\theta}E_{\theta}(x)]-\mathbb{E}_{x\sim p_{\theta}(x)}[-\beta(x)\nabla_{\theta}E_{\theta}(x)]$ (7)
where $\alpha(x)=1-\sigma(-E_{\theta}(x))$ and $\beta(x)=\sigma(-E_{\theta}(x))$ are data-dependent scaling factors, and $\sigma$ denotes the logistic sigmoid function. This formulation preserves the structural form of Equation˜5 while introducing adaptive scaling factors that modulate gradient contributions according to the model’s current energy assignments. According to Yin et al. (2022), these scaling factors stabilize training by providing automatic gradient regularization: as $E_{\theta}(x)$ increases for $p_{\text{data}}$ samples, the corresponding scaling factor $\alpha(x)=1-\sigma(-E_{\theta}(x))$ approaches zero, thereby attenuating the gradient contribution from such samples and preventing numerical overflow; conversely, when $E_{\theta}(x)$ becomes very negative for contrastive samples, $\beta(x)=\sigma(-E_{\theta}(x))$ approaches zero, preventing numerical underflow. In contrast, the standard EBM gradient (Equation˜5) is unconstrained and permits $E_{\theta}(x)$ to achieve arbitrarily large or small magnitudes, resulting in numerical instability during optimization. This gradient formulation stabilizes training at the cost of limiting the EBM to modeling the support of $p_{\text{data}}$ rather than learning the full density. We provide a formal characterization of the learned distribution in Section A.2, where we show that the optimal solution under our joint objective learns $f_{\theta}^{*}(x)[y]=\log p_{\text{data}}(y|x)$ on the support with constant marginal energy $E_{\theta}^{*}(x)=0$.
In addition to the above gradient reformualtion, the sampling required to estimate $\mathbb{E}_{x\sim p_{\theta}(x)}[\cdot]$ is performed using the PGD attack (Madry et al., 2017) instead of SGLD. Specifically, the contrastive samples $x$ from the model distribution are generated by initializing from an auxiliary out-of-distribution dataset $p_{\text{ood}}$ (e.g., the 80 million tiny images dataset for CIFAR-10 training) and performing $T$ iterations of gradient ascent on the negative energy function $E_{\theta}(x)$:
[…]
where $E_{\theta}(x)$ is the marginal energy function defined in Equation˜4, $\eta$ is the step size, and $T$ is the total number of PGD steps. Using the update direction suggested by Equation˜7 is equivalent to minimizing the Binary Cross-Entropy (BCE) loss:
$\mathcal{L}_{\text{BCE}}(\theta)=-\mathbb{E}_{x\sim p_{\text{data}}(x)}[\log(\sigma(-E_{\theta}(x)))]-\mathbb{E}_{x\sim p_{\theta}(x)}[\log(1-\sigma(-E_{\theta}(x)))]$ (9)
(i.e., $\nabla_{\theta}\mathcal{L}_{\text{BCE}}$ equals the right-hand side of Equation˜7). Minimizing this $\mathcal{L}_{\text{BCE}}$ implicitly trains the energy function $E_{\theta}(x)$ to assign low energy to data samples from $p_{\text{data}}(x)$ and high energy to the contrastive samples computed using the PGD attack.
We find this AT-based approach effectively addresses JEM’s training stability issues and produces high quality samples.
3.3 Classifier robustness and implicit regularization
Classifier robustness. While our AT-based approach improves the generative capabilities of JEM, the original JEM’s discriminative component still exhibits weak adversarial robustness compared to dedicated adversarially trained classifiers. To address this limitation, we complement our generative improvements by incorporating adversarial training for the discriminative term $p_{\theta}(y|x)$.
For each input sample $x$ with label $y$, we find an adversarial example $x_{adv}$ within an $\epsilon$-ball $B(x,\epsilon)$ around $x$ that maximizes the classification loss:
$x_{adv}=\operatorname*{arg\,max}_{x^{\prime}\in B(x,\epsilon)}\mathcal{L}_{\text{CE}}(\theta;x^{\prime},y)$ (10)
where $\mathcal{L}_{\text{CE}}(\theta;x^{\prime},y)$ is the standard cross-entropy loss and $B(x,\epsilon)$ is an $L_{p}$-norm ball. Similar to our generative component, we approximate this optimization using the PGD attack (Madry et al., 2017), generating adversarial examples through iterative gradient steps within the constraint set. The classification term is then defined as:
$\mathcal{L}_{\text{AT-CE}}(\theta)=\mathbb{E}_{(x,y)\sim p_{\text{data}}(x,y)}\left[-\log p_{\theta}(y|x_{adv})\right]$ (11)
Implicit regularization. Incorporating AT for the classifier not only ensures robust accuracy but also yields a synergistic benefit for the generative component. The original AT-EBMs framework (Yin et al., 2022) required explicit $R_{1}$ gradient penalties (Mescheder et al., 2018) for training stability. However, explicit gradient penalties directly constrain the model and introduce an additional hyperparameter to tune, while AT achieves similar regularization implicitly through adversarial data augmentation, potentially preserving more model expressiveness. Building on Roth et al. (2020), we show that AT implicitly bounds the $R_{1}$ penalty (Section A.3), eliminating the need for explicit gradient regularization. We empirically validate this: AT maintains bounded $R_{1}$ gradients throughout training, while standard training exhibits gradient explosion (Figure 2). This confirms that AT provides sufficient implicit regularization for stable EBM training without explicit gradient penalties.
3.4 Dual AT for joint modeling
Our complete model integrates adversarial training principles for both the generative and discriminative components, resulting in the combined objective:
$\mathcal{L}(\theta)=\mathcal{L}_{\text{AT-CE}}(\theta)+\mathcal{L}_{\text{BCE}}(\theta)$ (12)
where $\mathcal{L}_{\text{AT-CE}}(\theta)$ is the robust classification loss from Equation˜11, and $\mathcal{L}_{\text{BCE}}(\theta)$ is the AT-based generative loss from Equation˜9. This DAT approach simultaneously enhances the model’s discriminative robustness and generative capabilities, addressing the key limitations of the original JEM framework; full algorithmic details are provided in Appendix A.4.
Our approach shares conceptual similarities with RATIO (Augustin et al., 2020), which also combines adversarially robust classification with adversarial perturbations applied to out-of-distribution data:
$\mathcal{L}_{\text{RATIO}}(\theta)=\mathcal{L}_{\text{AT-CE}}(\theta)+\lambda\mathbb{E}_{x\sim p_{\text{ood}}(x)}\left[\max_{x^{\prime}\in B(x,\epsilon_{o})}\mathcal{L}_{\text{CE}}(\theta;x^{\prime},\mathbf{1}/K)\right]$ (13)
where $\mathbf{1}$ is the vector of all ones and $K$ is the number of classes. Despite this structural similarity, the approaches differ fundamentally in their objectives. RATIO’s secondary term attacks OOD samples to maximize classifier confidence, then penalizes this confidence via cross-entropy against a uniform distribution, explicitly targeting robust OOD detection. In contrast, our $\mathcal{L}_{\text{BCE}}(\theta)$ leverages AT-based energy function learning (Yin et al., 2022), using PGD to generate contrastive samples from OOD data and employing BCE loss to shape the energy landscape. While RATIO focuses primarily on reducing confidence in OOD regions, our approach prioritizes learning a stable and effective energy function that enables high-quality generative modeling alongside robust classification.
3.5 Two-stage training
Neural network architectures typically incorporate normalization layers to stabilize and speed up training: ResNet (He et al., 2016) uses batch normalization (BN) (Ioffe and Szegedy, 2015), while modern architectures like ConvNeXt (Liu et al., 2022) and Vision Transformers (Dosovitskiy et al., 2021; Vaswani et al., 2017) utilize layer normalization (Ba et al., 2016). Training energy-based joint models presents unique challenges with normalization layers. In particular, batch normalization has been identified as problematic for EBM training (Grathwohl et al., 2019; Yin et al., 2022; Zhao et al., 2020). Consistent with these findings, we observe that enabling BN during joint training destabilizes the optimization of the generative modeling term $\mathcal{L}_{\text{BCE}}$, leading to oscillating losses and failure to converge.
To address these challenges while maintaining the benefits of normalization during discriminative training, we propose a two-stage training strategy that generalizes effectively across diverse architectures:
  • Stage 1: Discriminative training. We first train the network with its original normalization configuration, optimizing only the robust classification objective $\mathcal{L}_{\text{AT-CE}}$ (Equation˜11). This stage is equivalent to standard adversarial training and leverages normalization layers to achieve faster convergence and strong robust classification performance. Notably, this stage can be skipped when pretrained robust classifiers are available, making our approach immediately applicable to existing robust models.
  • Stage 2: Joint training. After robust discriminative training, we modify the normalization behavior when necessary and continue training with the complete objective $\mathcal{L}(\theta)=\mathcal{L}_{\text{AT-CE}}(\theta)+\mathcal{L}_{\text{BCE}}(\theta)$ (Equation˜12). For architectures with batch normalization (ResNet, WRN), we disable BN by setting BN modules to eval mode, which freezes the BN statistics computed during Stage 1. For architectures with layer normalization (ConvNeXt), we maintain the normalization as-is.
This strategy not only addresses the incompatibility between batch normalization and EBM training, but also enables leveraging pretrained robust classifiers to dramatically reduce training costs (see Section˜B.2 for detailed computational analysis). As demonstrated in Section˜4.3, Stage 2 improves the generative modeling performance of pretrained robust classifiers while incurring minimal impact on the robust accuracy established in Stage 1 (see Section˜B.3 for training dynamics). Importantly, the two-stage training strategy works effectively for both ResNet and ConvNeXt models, making it suitable for architectures with better scaling properties such as Vision Transformers (Dosovitskiy et al., 2021; Singh et al., 2023; Peebles and Xie, 2023b).
3.6 Data augmentation
Strong data augmentations are necessary for achieving robust classification (Rebuffi et al., 2021; Gowal et al., 2020) but can distort the data distribution in ways detrimental to generative modeling. We therefore follow Yang et al. (2023) and apply separate augmentation strategies to the discriminative and generative components. While Yang et al. (2023) concludes that augmentations like random cropping with padding should be excluded from generative training to avoid artifacts like black borders, we find this is not a limitation in our framework. Notably, even with random cropping and padding applied, our generated samples do not inherit these artifacts, allowing us to improve robustness without degrading sample quality (see Section˜B.1.2). We therefore apply strong augmentations to the discriminative term $\mathcal{L}_{\text{AT-CE}}$ and mild augmentations to the generative term $\mathcal{L}_{\text{BCE}}$.
4 Experiments
4.1 Training setup
Datasets and architectures. We evaluate our approach on CIFAR-10 (Krizhevsky et al., 2009), CIFAR-100 (Krizhevsky et al., 2009), and ImageNet (Deng et al., 2009). For CIFAR-10/100 experiments, we use WRN-34-10 (Zagoruyko and Komodakis, 2016) following the official RATIO implementation. For ImageNet experiments, we employ ResNet-50 (He et al., 2016), WRN-50-4 (Zagoruyko and Komodakis, 2016), and ConvNeXt-Large with ConvStem (Singh et al., 2023).
Two-stage training. Since Stage 1 training is equivalent to standard adversarial training, we use pretrained standard AT checkpoints when available: a standard AT checkpoint from the RATIO codebase (Augustin et al., 2020) for CIFAR-10, pretrained ImageNet ResNet-50 and WRN-50-4 models from Salman et al. (2020), and pretrained ConvNeXt-Large with ConvStem from Singh et al. (2023) (originally trained for $\ell_{\infty}=4/255$ robustness), while training our own CIFAR-100 model following Augustin et al. (2020). For Stage 2 training, we initialize from the Stage 1 model and continue joint training. For ResNet and WRN architectures, we set the BN modules to eval mode (which disables BN while preserving the BN statistics computed during Stage 1). Complete training hyperparameters can be found in Section˜B.1.1.
Data augmentation. We employ separate data augmentation strategies for Stage 2 training: strong augmentations for $\mathcal{L}_{\text{AT-CE}}$ and basic transformations for $\mathcal{L}_{\text{BCE}}$ to preserve the data distribution. Detailed specifications can be found in Section˜B.1.2.
Out-of-distribution data. Same as RATIO, we use the 80 million tiny images (Torralba et al., 2008) as the OOD dataset ($p_{\text{ood}}$) for CIFAR-10/100 experiments. For ImageNet, as there are no established OOD datasets, we follow OpenImage-O (Wang et al., 2022a) and construct an OOD dataset from Open Images training set (Krasin et al., 2016). We randomly sample 350K images, restricting our selection to those whose labels do not overlap with any ImageNet classes, yielding 300K samples for training and 50K for FID evaluation.
4.2 Evaluation metrics
We measure both classification and generative modeling performance. For classification, we report clean accuracy and robust accuracy against $L_{2}$ attacks ($\epsilon=0.5$ for CIFAR-10/100 and $\epsilon=3.0$ for ImageNet) computed using AutoAttack (Croce and Hein, 2020). For generative modeling, we evaluate sample diversity and visual fidelity using Fréchet Inception Distance (FID) (Heusel et al., 2017) and Inception Score (IS) (Salimans et al., 2016). We focus on conditional generation; details of the generation setup are provided in Section˜B.4.
To measure the quality of counterfactuals, we generate sets of counterfactual examples by applying targeted attacks to training samples across a range of perturbation limits. For each target class, we compute the class-wise FID score between the set of counterfactuals targeted at that class and the set of training samples from the same class. Note that counterfactuals are generated by applying PGD attacks to in-distribution training samples, whereas generative modeling samples are created by applying PGD attacks to OOD inputs.
4.3 Results
We evaluate DAT on CIFAR-10, CIFAR-100, and ImageNet $\times256$ (Tables˜1 and 2). As detailed below, our approach achieves the first successful scaling of EBM-based hybrids to high resolutions while establishing a unique duality between robust classification and high-fidelity generation.
First EBM-based hybrid to scale to high-resolution datasets with adversarial robustness. Prior EBM-based hybrid approaches (JEM, SADA-JEM) are not explicitly optimized for adversarial robustness. On CIFAR-10, these methods achieve significantly lower robust accuracy than standard AT: JEM achieves 40.5% and SADA-JEM achieves 31.93%, compared to 75.73% for standard AT (Table˜1). Our approach addresses this limitation, achieving 75.75% robust accuracy—comparable to standard AT—while improving generative quality over prior EBM hybrids: FID 9.12 versus 38.4 (JEM) and 9.41 (SADA-JEM). Beyond robustness, prior EBM-based hybrids could not scale beyond low resolution or achieve competitive generative performance on ImageNet-level datasets. Our approach is the first to overcome this limitation: on ImageNet $\times256$, our ConvNeXt-Large model achieves FID 3.29 (Table˜2) with classification performance comparable to standard AT (Section˜C.6), demonstrating that EBM-based hybrid models can scale reliably to complex, high-resolution datasets while achieving strong classification performance.
Unique synergy of high-fidelity generation and adversarial robustness. While other scalable hybrids exist, they do not simultaneously achieve both strong adversarial robustness and state-of-the-art generative quality. The diffusion-based EGC achieves 13.56% robust accuracy on ImageNet compared to our 56.40%, while also achieving worse FID (6.05 vs. 3.29) (Table˜2). Similarly, RATIO targets robustness but not generation quality (FID 21.96 vs. our 9.12 on CIFAR-10). Qualitatively, Figures 6, 7, and 8 show that our method produces visually superior samples with fewer artifacts compared to RATIO and standard AT. This unique combination of generative quality and robustness enables critical applications such as robust counterfactual explanations (Section˜4.3.2), where our model produces substantially higher-quality counterfactuals than both non-robust and robustness-only methods.
[…]
4.3.1 Analysis and ablations
Generative-discriminative trade-off. Our experiments reveal that the number of PGD iterations $T$ (Equation˜8) controls the balance between discriminative and generative performance. On CIFAR-10, increasing $T$ from 40 to 50 improves FID from 9.12 to 7.57 at the cost of standard and robust accuracy. A similar trend is observed across all datasets. In Section˜C.7.3, we investigate this inherent tension, revealing that the generative and discriminative objectives encourage different representation alignments. We further show that this trade-off can be explicitly tuned through loss weighting in addition to PGD iterations.
Effect of model capacity and architecture. Our experiments on ImageNet demonstrate the benefits of increased model capacity and modern architectures. Scaling from ResNet-50 (26M parameters) to WRN-50-4 (223M parameters) yields consistent improvements across both discriminative and generative metrics. Beyond capacity, using state-of-the-art architectures also provides substantial benefits: ConvNeXt-L (198M parameters) substantially outperforms WRN-50-4 (223M parameters) in both accuracy and generation quality despite having fewer parameters, demonstrating the importance of architectural design alongside model scale.
Component contributions. We isolate the contributions of our key components, showing that the generative loss and decoupled augmentation are essential for high-fidelity synthesis (Section˜C.7.1).
OOD data efficiency. We also demonstrate high data efficiency, achieving strong performance even with limited auxiliary OOD data (Section˜C.7.2).
Robustness and stability. Beyond adversarial robustness, DAT maintains strong performance on common corruptions comparable to standard AT (Section˜C.5) and successfully generalizes to $L_{\infty}$ training (Section˜C.6). We also demonstrate high reproducibility with zero training divergence across all runs (Section˜C.8).
[…]
Table 1: Classification and generative modeling results on CIFAR-10 and CIFAR-100.
Method Acc% $\uparrow$ Robust Acc% $\uparrow$ IS $\uparrow$ FID $\downarrow$ CIFAR-10 hybrid models Residual Flow (Chen et al., 2019) 70.3 3.6 46.4 Glow (Kingma and Dhariwal, 2018) 67.6 3.92 48.9 IGEBM (Du and Mordatch, 2019) 49.1 8.3 37.9 JEM (Grathwohl et al., 2019) 92.9 40.5 8.76 38.4 VERA (Grathwohl et al., 2021) 93.2 8.11 30.5 JEM++ (Yang and Ji, 2021) 94.1 8.11 38.0 JEAT (Zhu et al., 2021) 85.16 8.80 38.24 Robust-JEM (Korst and Asadulaev, 2022) 8.71 41.17 SADA-JEM (Yang et al., 2023) 95.5 31.93 8.77 9.41 WEAT (Mirza et al., 2024) 83.36 8.97 30.74 EGC (Guo et al., 2023) 95.9 9.43 3.30 Joint-Diffusion (Deja et al., 2023) 96.4 6.400 RATIO (Augustin et al., 2020) 92.23 76.25 9.61 21.96 Standard AT (Augustin et al., 2020) 92.43 75.73 9.58 28.41 DAT ($T=40$) 91.92 75.75 9.92 9.12 DAT ($T=50$) 90.72 74.65 9.86 7.57 CIFAR-10 conditional generative models SNGAN (Miyato et al., 2018) 8.59 25.5 BigGAN (Brock et al., 2018) 9.22 14.73 StyleGAN2 (Karras et al., 2020b) 9.53 6.96 StyleGAN2 ADA (Karras et al., 2020a) 10.24 3.49 EDM (Karras et al., 2022) 1.79 CIFAR-100 hybrid models Joint-Diffusion (Deja et al., 2023) 77.6 16.8 SADA-JEM (Yang et al., 2023) 75.0 11.63 14.4 EGC (Guo et al., 2023) 77.9 11.50 4.88 RATIO (Augustin et al., 2020) 71.58 47.74 9.28 24.17 Standard AT (Augustin et al., 2020) 72.16 47.78 9.54 23.59 DAT ($T=45$) 65.76 45.94 10.99 10.73 DAT ($T=50$) 60.12 42.55 11.12 9.53
Table 2: Classification and conditional generative modeling results on ImageNet $\times256$.
Method Acc% $\uparrow$ Robust Acc% $\uparrow$ FID $\downarrow$ IS $\uparrow$ Params Steps Hybrid models EGC (Guo et al., 2023) 78.90 13.56 6.05 231.3 543M (U-Net) 1000 Standard AT (Salman et al., 2020) 64.91 39.96 15.12 286.2 26M (ResNet-50) 13 DAT ($T=15$) 61.31 39.96 6.87 322.65 26M (ResNet-50) 14 DAT ($T=30$) 55.96 37.14 5.28 319.3 26M (ResNet-50) 14 Standard AT (Salman et al., 2020) 71.25 45.86 37.33 260.2 223M (WRN-50-4) 12 DAT ($T=30$) 64.45 45.84 6.23 341.0 223M (WRN-50-4) 17 DAT ($T=65$) 58.78 40.74 4.94 358.0 223M (WRN-50-4) 19 Standard AT (Singh et al., 2023) 78.25 33.38 44.46 27.32 198M (ConvNext-L-CvSt) 0 DAT ($T=110$) 75.78 56.40 3.29 310.2 198M (ConvNext-L-CvSt) 36 Conditional generative models BigGAN-deep (Brock et al., 2018) 6.95 203.6 340M (ResNet) 1 ADM-G (Dhariwal and Nichol, 2021) 4.59 186.7 608M (U-Net) 250 LDM-4-G (Rombach et al., 2022) 3.60 247.7 400M (U-Net) 250 DiT-XL/2-G (Peebles and Xie, 2023a) 2.27 278.2 675M (Transformers) 250 VAR-d16 (Tian et al., 2024) 3.30 274.4 310M (Transformers) 10 VAR-d30-re (Tian et al., 2024) 1.73 350.2 2.0B (Transformers) 10
4.3.2 Counterfactual generation, OOD detection, and calibration
Counterfactual generation. Visual counterfactual explanations (VCEs) reveal what minimal semantic changes would flip a classifier’s decision. Standard classifiers cannot generate meaningful VCEs because their input gradients lack semantic structure—gradient-based modifications produce adversarial noise rather than interpretable features. Prior VCE methods were therefore restricted to adversarially robust models (Augustin et al., 2020; Boreiko et al., 2022). DVCE (Augustin et al., 2022) overcomes this limitation by guiding an unconditional diffusion model using gradients from an auxiliary adversarially robust classifier, projecting these onto a cone around the target classifier’s gradient. This enables VCEs for arbitrary classifiers but requires multiple external dependencies: a pretrained diffusion model, a robust classifier, and careful hyperparameter tuning.
In contrast, DAT is intrinsically explainable: it generates VCEs directly by gradient descent on the joint energy $E_{\theta}(x,y)$, requiring no external models. Because our joint objective explicitly learns the energy landscape of the joint distribution $p(x,y)$, the resulting gradients $\nabla_{x}E_{\theta}(x,y)$ naturally point toward semantically valid configurations. Therefore, our model’s improved generative capability directly translates to higher-quality counterfactual explanations.
Figure˜1 compares counterfactual quality across different models while accounting for classifier confidence. Our approach consistently generates counterfactuals with lower FIDs than baseline methods when achieving similar target class confidence. For instance, when the RATIO baseline reaches approximately 0.89 confidence in the target class (at $\epsilon=8$), its corresponding FID is 43.18. Our DAT model achieves a similar confidence level at $\epsilon=4$ with a significantly better FID of 25.53. This demonstrates that, for a comparable level of certainty that the counterfactual represents the target class, our generated samples are substantially more faithful to the true visual characteristics of that class, indicating more plausible counterfactuals. We provide visualizations of counterfactuals in Section˜C.2.
Figure 1: Counterfactual FIDs and classifier confidences under different perturbations.
OOD detection. Our approach generally underperforms RATIO on OOD detection. Ablation studies show this gap persists even when using identical aggressive augmentation for both the generative and discriminative components, indicating it stems from fundamental objective differences rather than the use of milder augmentation for the generative term: RATIO explicitly optimizes for OOD detection while our generative loss prioritizes learning accurate energy functions for generation. The complete details can be found in Section˜C.3
Calibration. Our model’s calibration performance is dataset-dependent, with detailed results provided in Section˜C.4. While the model is well-calibrated on CIFAR-10, outperforming the standard AT and RATIO baselines, it exhibits higher overconfidence on CIFAR-100 and ImageNet. The results suggest that prioritizing generative quality may come at the cost of calibration.
5 Conclusion
We presented Dual Adversarial Training (DAT), demonstrating that hybrid models can achieve state-of-the-art performance without compromise. Our approach advances hybrid modeling in three ways: (1) we present the first EBM-based hybrid to scale to high-resolution complex datasets, achieving FID 3.29 and competitive robust accuracy on ImageNet $\times256$; (2) we uniquely combine state-of-the-art generative quality with adversarial robustness, enabling critical applications such as robust counterfactual explanations; and (3) as a standalone generative model, our approach matches the generative quality of state-of-the-art autoregressive methods and surpasses leading diffusion models while offering unique versatility for diverse synthesis tasks.
Future work could advance this approach in several directions: improving training efficiency with persistent markov chains, scaling the framework to higher-capacity architectures such as ConvNeXt-XLarge and Vision Transformers; improving secondary tasks like out-of-distribution detection by developing hybrid objectives that combine our generative loss with RATIO’s; and applying the model to broader image synthesis tasks such as those demonstrated by Santurkar et al. (2019).
Appendix A Theory & Method
A.1 Extended related work
Connections between adversarial robustness and energy-based models Zhu et al. [2021] reinterpret adversarially trained classifiers as joint energy-based models, showing that adversarial training implicitly flattens the energy landscape around real data by reducing the energy of nearby high-energy adversarial examples. They identify $E_{\theta}(x,y)$ as the key energy term for conditional generation and propose JEAT, which employs energy-based adversarial perturbations and SGLD sampling for likelihood estimation and generation. Mirza et al. [2024] extend this analysis by decomposing the cross-entropy loss as $\mathcal{L}_{\text{CE}}(x,y;\theta)=E_{\theta}(x,y)-E_{\theta}(x)$, revealing that untargeted attacks increase the joint energy $E_{\theta}(x^{*},y)$ (reducing classifier confidence of class $y$) while decreasing the marginal energy $E_{\theta}(x^{*})$. They show that robust overfitting corresponds to divergence between $E_{\theta}(x)$ and $E_{\theta}(x^{*})$, and that state-of-the-art robust models achieve better generalization by smoothing the marginal energy landscape around natural data. Wang et al. [2022b] proposed a unified Contrastive Energy-based Model (CEM) framework that interprets adversarial training as biased maximum likelihood estimation of an energy-based model $p_{\theta}(x,y)=\exp(f_{\theta}(x,y))/Z(\theta)$. Unlike JEM [Grathwohl et al., 2019] which samples negative examples from random noise via Langevin dynamics, CEM shows that PGD-generated adversarial perturbations from real data serve as implicit negative samples, providing more stable training without requiring random noise or OOD data. The framework unifies supervised (P-CEM) and unsupervised (NP-CEM) scenarios, revealing connections between adversarial training, contrastive learning, and energy-based modeling, and enables improved sampling algorithms that achieve state-of-the-art generative performance. While these works provide valuable insights into the energy-based interpretation of adversarial training, training methods that rely on SGLD-based sampling (e.g., JEAT) inherit the instabilities of standard JEM [Grathwohl et al., 2019]. In contrast, our work addresses these stability issues by employing PGD-based contrastive sampling and replacing the unstable EBM loss with a standard BCE loss.
Learning EBMs with adversarial training Yin et al. [2022] explored an alternative approach to learning EBMs by leveraging the mechanism of Adversarial Training (AT). They established a connection between the objective of binary AT (discriminating real data from adversarially perturbed out-of-distribution data) and the SGLD-based maximum likelihood training commonly used for EBMs. Specifically, they showed that the binary classifier learned via AT implicitly defines an energy function that models the support of the data distribution, assigning low energy to in-distribution regions and high energy to out-of-distribution (OOD) regions. The PGD attack used in AT to generate adversarial samples from OOD data was interpreted as a non-convergent sampler that produces contrastive data, analogous to MCMC sampling in EBM training. Although the resulting energy function can only capture the support rather than recover the exact density, their model achieves competitive image generation performance compared to explicit EBMs. Notably, this AT-based EBM learning approach is more stable than traditional MCMC-based EBM training and demonstrated strong performance in worst-case out-of-distribution detection, similar to methods like RATIO [Augustin et al., 2020]. However, AT-EBM focuses on unconditional generative modeling and employs an explicit $R_{1}$ gradient penalty to stabilize training, which can constrain model expressiveness. Our work incorporates AT-based EBM learning into the JEM framework to perform conditional generative modeling with implicit $R_{1}$ regularization from adversarial training, using ancestral sampling from the conditional distribution $p(x|y)$ rather than the marginal distribution $p(x)$.
Improving joint energy-based models Building on the original JEM framework [Grathwohl et al., 2019], several works have explored techniques to improve training stability and performance. Yang and Ji [2021] (JEM++) introduced multiple training improvements: (1) Proximal SGLD that constrains samples within an $L_{p}$-norm ball of previous samples via gradient clamping for improved stability; (2) YOPO-inspired acceleration (PYLD) that reduces redundant backpropagation by exploiting the coupling between samples and first-layer weights; (3) Informative initialization from a class-conditional Gaussian mixture distribution estimated from training data, which accelerates SGLD convergence, improves stability, and enables batch normalization. Korst and Asadulaev [2022] (Robust-JEM) further enhanced JEM++ by incorporating adversarial training into the discriminative component, empirically observing improved training stability. At inference time, they propose a "combined inference" approach where initial samples from PGD adversarial attacks are refined using SGLD, improving generative performance. However, both JEM++ and Robust-JEM fundamentally still rely on SGLD for sampling and MLE-based objectives for the generative component, inheriting SGLD’s intrinsic instability issues. While these works introduced valuable techniques for improving SGLD stability, they did not fundamentally resolve the instability of SGLD-based training and both remain limited to CIFAR-scale ($\times32$) datasets. Our approach departs fundamentally from this line of work by: (1) providing mathematical analysis and empirical evidence showing that adversarial training offers implicit $R_{1}$ regularization (Section A.3); (2) replacing the MLE-based objective with BCE-based gradients; (3) using deterministic PGD-based sampling instead of stochastic Langevin dynamics; and (4) introducing a two-stage training strategy to address the incompatibility between batch normalization and EBM training. This enables scaling to high-resolution ImageNet synthesis ($\times256$) with state-of-the-art generative quality.
In- and out-distribution adversarial robustness Addressing the multifaceted challenge of creating models that are simultaneously accurate, robust, and reliable on out-of-distribution (OOD) data, Augustin et al. [2020] proposed RATIO (Robustness via Adversarial Training on In- and Out-distribution). Their approach combines standard adversarial training (AT) on the in-distribution data, aimed at improving robustness against adversarial examples, with a form of AT on OOD data, which enforces low and uniform confidence predictions within a neighborhood around OOD samples. The combined objective trains the model to maintain correct, robust classifications for in-distribution data while actively discouraging high-confidence predictions for OOD inputs, even under adversarial manipulation. Augustin et al. [2020] demonstrated that RATIO achieves state-of-the-art $L_{2}$ robustness on datasets like CIFAR-10, often with less degradation in clean accuracy compared to standard AT alone. Furthermore, they showed that RATIO yields reliable OOD detection performance, particularly in worst-case scenarios where OOD samples are adversarially perturbed to maximize confidence. Their work also highlighted that the $L_{2}$ robustness fostered by RATIO enables the generation of meaningful visual counterfactual explanations directly in pixel space, where optimizing confidence towards a target class results in the emergence of corresponding class-specific visual features.
Robust classifiers for image synthesis and manipulation Santurkar et al. [2019] demonstrated that adversarially robust classifiers can serve as powerful primitives for diverse image synthesis tasks. The core insight of their work is that the process of adversarial training—which optimizes the worst-case loss over an $\ell_{2}$ perturbation set rather than expected loss—compels a model to learn more perceptually aligned and human-interpretable feature representations by preventing reliance on imperceptible artifacts. Based on this insight, Santurkar et al. [2019] showed that simple gradient ascent on class scores from such robust classifiers enables a unified framework for image generation, inpainting, image-to-image translation, super-resolution, and interactive manipulation—tasks typically requiring specialized GAN architectures or complex generative models.
A.2 Theoretical characterization
We provide a formal characterization of what distribution our joint objective (Equation˜9) learns. Our analysis builds on the theoretical framework from Yin et al. [2022] (AT-EBM) and extends it to the conditional modeling setting by deriving the optimal class logits under the joint objective.
Optimal solution for class logits $f_{\theta}^{*}(x)[y]$.
Following Yin et al. [2022], the generative component of our objective can be expressed as a maximin optimization problem. Let $D(x)=\sigma(-E_{\theta}(x))$ where $E_{\theta}(x)=-\log\sum_{y}\exp(f_{\theta}(x)[y])$ is the marginal energy function. For the theoretical analysis below, we assume (1) the PGD attack in Equation˜8 converges to the global minimum of $E_{\theta}(x)$, (2) the model has sufficient capacity, and (3) infinite training data from $p_{\text{data}}(x,y)$. Under these assumptions, minimizing $\mathcal{L}_{\text{BCE}}(\theta)$ implicitly solves:
$\max_{D}\min_{p_{T}}U(D,p_{T})=\mathbb{E}_{x\sim p_{\text{data}}(x)}[\log D(x)]+\mathbb{E}_{x\sim p_{T}}[\log(1-D(x))]$ (14)
where $p_{T}$ represents the distribution of samples after PGD attack initialized from the auxiliary out-of-distribution dataset $p_{\text{ood}}$.
Under these assumptions, the optimal solution to Equation˜14 is characterized by Proposition 1 of Yin et al. [2022], which shows that at optimum $U(D^{*},p_{T}^{*})=-\log(4)$ with:
  1. 1.
    $D^{*}(x)=\frac{1}{2}$ for all $x\in\mathrm{Supp}(p_{\text{data}}(x))$
  2. 2.
    $D^{*}(x)\leq\frac{1}{2}$ for all $x\notin\mathrm{Supp}(p_{\text{data}}(x))$
  3. 3.
    $p_{T}^{*}$ is supported on $\{x:D(x)=\frac{1}{2}\}$
where $\mathrm{Supp}(p_{\text{data}}(x))$ denotes the support of the marginal data distribution.
The above result characterizes only the marginal energy $E_{\theta}(x)$, leaving the individual class logits $f_{\theta}(x)[y]$ underdetermined. We now derive their optimal values by incorporating the discriminative objective $\mathcal{L}_{\text{AT-CE}}$.
Proposition A.1 (Optimal class logits).
Under the assumptions stated above, at the optimal solution to the joint objective $\mathcal{L}(\theta)=\mathcal{L}_{\text{AT-CE}}(\theta)+\mathcal{L}_{\text{BCE}}(\theta)$, the class logits satisfy on the support:
$f_{\theta}^{*}(x)[y]=\log p_{\text{data}}(y|x)\quad\text{for all }x\in\mathrm{Supp}(p_{\text{data}}(x))$ (15)
Proof.
From the AT-EBM result above, on the support we have $D^{*}(x)=\sigma(-E_{\theta}^{*}(x))=\frac{1}{2}$, which implies:
$E_{\theta}^{*}(x)=0\implies-\log\sum_{y}\exp(f_{\theta}^{*}(x)[y])=0\implies\sum_{y}\exp(f_{\theta}^{*}(x)[y])=1$ (16)
The conditional distribution is defined as:
$p_{\theta}(y|x)=\frac{\exp(f_{\theta}(x)[y])}{\sum_{y^{\prime}}\exp(f_{\theta}(x)[y^{\prime}])}$ (17)
Substituting the constraint from Equation˜16:
$p_{\theta}^{*}(y|x)=\frac{\exp(f_{\theta}^{*}(x)[y])}{1}=\exp(f_{\theta}^{*}(x)[y])$ (18)
The adversarial cross-entropy objective $\mathcal{L}_{\text{AT-CE}}$ minimizes $\log p_{\theta}(y|x)$ over worst-case perturbations of $(x,y)\sim p_{\text{data}}(x,y)$. Under our capacity assumption, at optimality on the support this reduces to standard cross-entropy behavior, yielding $p_{\theta}^{*}(y|x)=p_{\text{data}}(y|x)$. Since the optimal predictive distribution naturally satisfies the normalization constraint $\sum_{y}p_{\text{data}}(y|x)=1$, the solution to the discriminative objective is fully compatible with the marginal energy constraint $\sum_{y}\exp(f_{\theta}^{*}(x)[y])=1$ derived from the generative objective. Therefore:
$\exp(f_{\theta}^{*}(x)[y])=p_{\text{data}}(y|x)$ (19)
Taking logarithms gives Equation˜15. ∎
On-support behavior.
Proposition A.1 implies that on the support: (1) the joint energy equals the negative conditional log-probability $E_{\theta}^{*}(x,y)=-f_{\theta}^{*}(x)[y]=-\log p_{\text{data}}(y|x)$, (2) the marginal energy is constant $E_{\theta}^{*}(x)=0$, and (3) the unnormalized marginal density is constant, $p_{\theta}^{*}(x)\propto\exp(-E_{\theta}^{*}(x))=1$. This confirms that the model assigns equal density to all points on the support, rather than learning the true data density $p_{\text{data}}(x)$. For datasets with deterministic labels where $p_{\text{data}}(y|x)=\delta_{y,y_{\text{true}}(x)}$, this implies $f_{\theta}^{*}(x)[y]=0$ for the true class and $f_{\theta}^{*}(x)[y]=$-$\infty$ otherwise.
Off-support behavior.
For $x\notin\mathrm{Supp}(p_{\text{data}}(x))$, the optimal solution to Equation˜14 constrains $D^{*}(x)\leq\frac{1}{2}$, which implies $E_{\theta}^{*}(x)\geq 0$ and thus $\sum_{y}\exp(f_{\theta}^{*}(x)[y])\leq 1$. Since $\mathcal{L}_{\text{AT-CE}}$ is only computed on the data support and its adversarial perturbations, the individual class logits are underdetermined for points far from the support. From the constraint $\sum_{y}\exp(f_{\theta}^{*}(x)[y])\leq 1$, we have $f_{\theta}^{*}(x)[y]\leq 0$ for all classes $y$, and consequently $E_{\theta}^{*}(x,y)=-f_{\theta}^{*}(x)[y]\geq 0$.
Comparison.
For datasets with deterministic labels, this creates a hierarchical energy structure:
  • Valid pairs $(x\in\mathrm{Supp},y=y_{\text{true}})$: The joint energy is exactly $E_{\theta}^{*}(x,y)=0$.
  • On-support, incorrect labels $(x\in\mathrm{Supp},y\neq y_{\text{true}})$: The joint energy diverges to $E_{\theta}^{*}(x,y)=\infty$.
  • Off-support (OOD) $(x\notin\mathrm{Supp},\text{any }y)$: The joint energy is bounded below, $E_{\theta}^{*}(x,y)\geq 0$, but the exact value is underdetermined.
Thus, the learned energy function $E_{\theta}(x,y)$ enables robust classification (via $\arg\min_{y}E_{\theta}(x,y)$), generation (via minimizing $E_{\theta}(x,y)$ over $x$), and OOD detection (via thresholding $\min_{y}E_{\theta}(x,y)$).
Finite-step vs. convergent PGD.
The theoretical analysis above assumes the PGD attack converges to the global minimum of $E_{\theta}(x)$, which leads to the maximin formulation in Equation˜14. In practice, however, we use finite-step PGD (Equation˜8), which does not explicitly solve the inner minimization problem, creating a crucial gap between theory and practice. Our experiments demonstrate that our approach achieves diverse, high-quality generation, suggesting that rather than converging to energy minima, finite-step PGD behaves like a sampler in practice: starting from different initializations in the OOD dataset, it explores different regions of the energy landscape, leading to diverse generated samples.
Summary and implications.
Our formal analysis reveals that our joint objective learns a fundamentally different quantity than MLE-based methods. While MLE-based JEM theoretically learns the full density $p_{\theta}(x)$ with a valid partition function (though in practice short-run SGLD fails to achieve this), our approach explicitly learns the support of the data distribution with $E_{\theta}(x)=0$ on the support. The optimal class logits $f_{\theta}^{*}(x)[y]=\log p_{\text{data}}(y|x)$ (Proposition A.1) reveal how the joint objective uniquely determines the solution on the support.
This support-based characterization has important implications. The constant marginal energy $E_{\theta}^{*}(x)=0$ on the support means the model assigns equal unnormalized density to all points on the support, theoretically discarding frequency information about the data distribution. This represents a significant theoretical limitation: the model cannot distinguish between common and rare examples within the support, and thus cannot perform density estimation tasks that require modeling relative frequencies.
Despite this limitation, the support-based formulation provides clear advantages: superior training stability compared to SGLD-based methods, robust classification, and effective OOD detection—benefits that do not require density information. Additionally, our strong empirical generation results demonstrate that the finite-step PGD dynamics discussed above act as an effective sampler, capturing density variations despite the theoretical prediction of uniformity. Overall, the support-based approach provides a practical and stable framework for joint modeling, trading full density estimation for superior training stability and strong empirical performance.
A.3 Implicit $R_{1}$ regularization
Building on the operator norm equivalence established by Roth et al. [2020], we show that adversarial training implicitly bounds the $R_{1}$ gradient penalty, eliminating the need for explicit gradient regularization. Consider a classifier $f:\mathbb{R}^{d}\to\mathbb{R}^{K}$ producing logits for $K$ classes.
$R_{1}$ regularization directly penalizes large gradients of the true class logit:
$\mathcal{L}_{R_{1}}=\mathbb{E}_{(x,y)\sim p_{\text{data}}}\left[\left\|\nabla_{x}f_{y}(x)\right\|_{2}^{2}\right]$ (20)
In practice, adversarial training uses cross-entropy loss to enforce consistent predictions:
$\mathcal{L}_{\text{AT}}=\mathbb{E}_{(x,y)\sim p_{\text{data}}}\left[\max_{\|\delta\|_{p}\leq\epsilon}L(f(x+\delta),y)\right]$ (21)
Adversarial training as operator norm regularization.
Roth et al. [2020] proved that $\ell_{p}$-norm constrained adversarial training with an $\ell_{q}$-norm loss on the logits is equivalent to data-dependent $(p,q)$-operator norm regularization:
$\mathcal{L}_{\text{AT}}\equiv\mathcal{L}_{\text{standard}}+\lambda(\epsilon)\cdot\|J_{f}(x)\|_{p,q}$ (22)
where $\lambda(\epsilon)$ is an implicit regularization coefficient that scales linearly with the attack budget $\epsilon$. Here $J_{f}(x)\in\mathbb{R}^{K\times d}$ denotes the input-Jacobian of the logits, and the $(p,q)$-operator norm is defined as:
$\|J_{f}(x)\|_{p,q}:=\max_{\|v\|_{p}=1}\|J_{f}(x)v\|_{q}$ (23)
This measures the maximal signal amplification when propagating a norm-bounded input perturbation through the linearized network. For $\ell_{2}$-constrained adversarial training, the $(2,2)$-operator norm reduces to the spectral norm $\sigma_{\max}(J_{f}(x))$, the largest singular value of the Jacobian:
$\mathcal{L}_{\text{AT}}\equiv\mathcal{L}_{\text{standard}}+\lambda(\epsilon)\cdot\sigma_{\max}(J_{f}(x))$ (24)
The optimal perturbation direction aligns with the dominant right singular vector of $J_{f}(x)$.
Connection to $R_{1}$ regularization.
Since adversarial training regularizes $\sigma_{\max}(J_{f}(x))$, we trace how this bound propagates to $\|\nabla_{x}f_{y}(x)\|_{2}$, which $R_{1}$ penalizes. By definition, the spectral norm satisfies $\sigma_{\max}(J_{f}(x))=\sigma_{\max}(J_{f}(x)^{T})=\max_{\|v\|_{2}=1}\|J_{f}(x)^{T}v\|_{2}$. Let $e_{y}\in\mathbb{R}^{K}$ be the standard basis vector for class $y$. The product $J_{f}(x)^{T}e_{y}$ extracts the $y$-th column of $J_{f}(x)^{T}$, which corresponds to the gradient $\nabla_{x}f_{y}(x)$. Since $\|e_{y}\|_{2}=1$, we have:
$\|\nabla_{x}f_{y}(x)\|_{2}=\|J_{f}(x)^{T}e_{y}\|_{2}\leq\max_{\|v\|_{2}=1}\|J_{f}(x)^{T}v\|_{2}=\sigma_{\max}(J_{f}(x))$ (25)
Thus, adversarial training’s regularization of $\sigma_{\max}(J_{f}(x))$ upper-bounds the $R_{1}$ gradient penalty $\|\nabla_{x}f_{y}(x)\|_{2}$.
Bound on all per-class gradients.
This analysis reveals that adversarial training provides a stronger regularization than $R_{1}$ alone. While $R_{1}$ penalizes only $\|\nabla_{x}f_{y}(x)\|_{2}$, adversarial training simultaneously bounds all per-class gradient norms:
$\|\nabla_{x}f_{k}(x)\|_{2}\leq\sigma_{\max}(J_{f}(x)),\quad\forall k\in\{1,\ldots,K\}$ (26)
This includes the probability-weighted mean gradient $\sum_{k}p_{k}\nabla_{x}f_{k}(x)$, since its norm is bounded by $\max_{k}\|\nabla_{x}f_{k}(x)\|_{2}$.
Effective regularization strength scales with attack budget.
An important consequence of this framework is that the effective regularization strength scales with the adversarial attack budget $\epsilon$. From the equivalence established above, the adversarial training objective implicitly regularizes $\sigma_{\max}(J_{f}(x))$ with strength $\lambda(\epsilon)$. For the network to maintain bounded total loss, the regularization term must remain bounded: $\lambda(\epsilon)\cdot\sigma_{\max}(J_{f}(x))\lesssim C$ for some constant $C$. Since $\lambda(\epsilon)$ scales linearly with $\epsilon$, this implies:
$\sigma_{\max}(J_{f}(x))\lesssim\frac{C}{\lambda(\epsilon)}$ (27)
Combining with our direct bound:
$\|\nabla_{x}f_{y}(x)\|_{2}\leq\sigma_{\max}(J_{f}(x))\lesssim\frac{C}{\lambda(\epsilon)}$ (28)
Thus, the $R_{1}$ gradient norm scales inversely with the regularization strength. Larger attack budgets $\epsilon$ induce stronger implicit regularization $\lambda(\epsilon)$, leading to lower $R_{1}$ gradients.
Connection to GAN regularization.
GAN training commonly employs spectral normalization [Miyato et al., 2018] on weight matrices and $R_{1}$ gradient penalties [Mescheder et al., 2018] for stability. The weight spectral norm provides a global upper bound on the Jacobian spectral norm: $\sigma(J_{f}(x))\leq\prod_{\ell}\sigma(W^{\ell})$ [Yoshida and Miyato, 2017]. However, Roth et al. [2020] show this bound can be loose as it ignores data-dependent activation patterns. AT’s equivalence to data-dependent spectral norm regularization is more targeted, directly controlling sensitivity at actual data points. Combined with our analysis showing this bounds $R_{1}$, AT provides a principled implicit alternative to both explicit GAN regularizers.
Empirical validation.
We validate this theoretical analysis by tracking $R_{1}$ gradient norms during Stage 2 joint training on ImageNet $\times256$ (ResNet-50, $T=15$). Figure 2 compares two settings: (1) adversarial training on the discriminative loss ($\mathcal{L}_{\text{AT-CE}}$), and (2) standard training on the discriminative loss. As predicted by the analysis above, the adversarial training curve remains bounded and stable throughout training, while standard training exhibits gradient explosion—with $R_{1}$ values increasing by orders of magnitude in later stages. While the exact equivalence from Roth et al. [2020] holds under technical conditions (small $\epsilon$ relative to ReLU cell size, infinite attack step size limit), our empirical observations are consistent with the theoretical predictions. This confirms that implicit regularization from adversarial training suffices to maintain bounded $R_{1}$ gradients, thereby enabling stable energy-based model training.
Figure 2: $R_{1}$ gradient norm during Stage 2 joint training on ImageNet $\times256$ (ResNet-50, $T=15$). Y-axis is in log scale. Adversarial training maintains bounded gradients while standard training exhibits gradient explosion.
A.4 Training algorithm
The complete training procedure for our combined objective (Equation˜12) is detailed in Algorithm˜1. We note that to train the generative component $\mathcal{L}_{\text{BCE}}$, we sample from $p_{\theta}(x)$ to estimate $\mathbb{E}_{x\sim p_{\theta}(x)}[-\nabla_{\theta}E_{\theta}(x)]$ in Equation˜5. In the context of JEM, there are broadly two strategies for drawing samples from $p_{\theta}(x)$ [Grathwohl et al., 2019]:
  1. 1.
    Direct sampling from the marginal distribution using gradient-based MCMC (e.g., SGLD or PGD) on the marginal energy $E_{\theta}(x)=-\log\sum_{y}\exp(f_{\theta}(x)[y])$, as implied by Equation˜8.
  2. 2.
    Ancestral sampling, which first draws a label $y\sim p_{\text{data}}(y)$, then samples $x\sim p_{\theta}(x|y)$ by running gradient-based MCMC on the joint energy $E_{\theta}(x,y)=-f_{\theta}(x)[y]$.
Although both approaches yield unbiased estimates, we find ancestral sampling to be practically superior for training stability, possibly because it leverages the classifier’s existing strong class representations to provide better mode coverage and mixing properties, while direct sampling from the marginal distribution often diverges. We also find ancestral sampling (conditional generation) yields substantially better FID than directly sampling from marginal distribution (see Table˜9).
Consequently, our implementation adopts ancestral sampling when generating contrastive samples (Algorithm˜1). Specifically, we first sample a label $y^{\prime}\sim p_{\text{data}}(y)$, then generate a contrastive sample $x_{T}$ by performing $T$ iterations of PGD on the negative joint energy function $E_{\theta}(x,y^{\prime})$, starting from an initial sample $x_{0}\sim p_{\text{ood}}$. This class-conditional contrastive sample $x_{T}$ is then used in the $\mathcal{L}_{\text{BCE}}$ objective (Equation˜9), whose gradient (Equation˜7) provide an approximation to Equation˜5.
Algorithm 1 DAT training: Given network logits $f_{\theta}$, in-distribution dataset $p_{\text{data}}$, auxiliary out-of-distribution dataset $p_{\text{ood}}$, classification AT bound $\epsilon$, PGD iterations $T$, PGD step size $\eta$
1:while not converged do
2:  Sample $(x,y)\sim p_{\text{data}}(x,y)$, apply aggressive augmentation to $x$
3:  Sample $\hat{x}\sim p_{\text{data}}(x)$, $x_{0}\sim p_{\text{ood}}(x)$, apply mild augmentation to $\hat{x}$ and $x_{0}$
4:  Solve $x_{adv}=\operatorname*{arg\,max}_{x^{\prime}\in B(x,\epsilon)}\mathcal{L}_{\text{CE}}(\theta;x^{\prime},y)$ via PGD attack
5:  $\mathcal{L}_{\text{AT-CE}}(\theta)=-\log p_{\theta}(y|x_{adv})$ $\triangleright$ Robust classification loss
6:  Initialize $x_{t}\leftarrow x_{0}$ for $t=0$, sample $y^{\prime}\sim p_{\text{data}}(y)$
7:  for $t\in\{1,\ldots,T\}$ do $\triangleright$ Generate contrastive sample for EBM
8:   $g=\nabla_{x}(-E_{\theta}(x_{t-1},y^{\prime}))$ $\triangleright$ Gradient of negative energy
[…]
10:  end for
11:  $\mathcal{L}_{\text{BCE}}(\theta)=-\log(\sigma(-E_{\theta}(\hat{x})))-\log(1-\sigma(-E_{\theta}(x_{T})))$ $\triangleright$ Generative modeling loss
12:  $\mathcal{L}(\theta)=\mathcal{L}_{\text{AT-CE}}(\theta)+\mathcal{L}_{\text{BCE}}(\theta)$
13:  Compute parameter gradients $\nabla_{\theta}\mathcal{L}(\theta)$ and update $\theta$
14:end while
Appendix B Training & Evaluation
B.1 Implementation details
B.1.1 Hyperparameters
We implement the two-stage training approach as described in Section 3.5. Table 3 summarizes the key hyperparameters used for both stages across different datasets.
For Stage 1, we utilize a pretrained CIFAR-10 model from RATIO [Augustin et al., 2020] and pretrained ImageNet ResNet-50 and WRN-50-4 models from Salman et al. [2020], while training our own CIFAR-100 model following the RATIO methodology with the hyperparameters specified in Table 3. For ConvNeXt-Large experiments on ImageNet, we use the pretrained robust checkpoint from Singh et al. [2023]. We select the EMA model with the best robust test accuracy as the final Stage 1 model.
For Stage 2, we initialize from the Stage 1 model and continue training. For ResNet and WRN architectures, we disable batch normalization by setting all BN modules to evaluation mode. During this stage, we optimize the complete objective function $\mathcal{L}(\theta)=\mathcal{L}_{\text{AT-CE}}(\theta)+\mathcal{L}_{\text{BCE}}(\theta)$ using fixed learning rates as specified in Table˜3. For the discriminative component $\mathcal{L}_{\text{AT-CE}}(\theta)$, CIFAR and ImageNet ResNet/WRN models continue to use the same adversarial settings as stage 1, while ConvNeXt transitions from $L_{\infty}$ to $L_{2}$ perturbations with adjusted PGD parameters (see Table˜4). The generative component $\mathcal{L}_{\text{BCE}}(\theta)$ employs the parameters detailed in Table˜5.
We select the Stage 2 checkpoint with the best FID score for the final evaluation reported in Section˜4.3.
Table 3: Training hyperparameters for both stages. Epochs for stage 2 are estimated based on the number of in-distribution images seen by the discriminative component during training.
CIFAR-10/100 ImageNet ImageNet Architecture WRN-34-10 ResNet-50/WRN-50-4 ConvNeXt-L-CvSt BatchNorm Enabled (stage 1), Disabled (stage 2) Enabled (stage 1), Disabled (stage 2) N/A LayerNorm N/A N/A Enabled Optimizer SGD with Nesterov SGD with Nesterov AdamW Weight decay $5\times 10^{-4}$ $1\times 10^{-4}$ (stage 1), $5\times 10^{-4}$ (stage 2) 0.05 Batch size 128 512 756 (stage 1), 512 (stage 2) EMA Enabled Enabled Enabled LR (stage 1) 0.1 (cosine schedule) 0.1 (step decay at epochs 30, 60, 90) 0.001 (cosine decay with warm-up) LR (stage 2) 0.001 (CIFAR-10), 0.009 (CIFAR-100) 0.001 0.0003 Epochs (stage 1) 300 90 100 Epochs (stage 2) 26 (CIFAR-10), 30 (CIFAR-100) 0.78 (ResNet-50), 0.39 (WRN-50-4) 1.09
Table 4: Adversarial training parameters for $\mathcal{L}_{\text{AT-CE}}$ (identical across stages for CIFAR and ImageNet ResNet/WRN).
CIFAR-10/100 ImageNet (ResNet/WRN) ImageNet (ConvNeXt-L-CvSt) PGD steps 10 2 2 PGD step size 0.1 2.0 2/255 ($L_{\infty}$, stage 1), 2.0 ($L_{2}$, stage 2) Perturbation bound $L_{2}$, $\epsilon=0.5$ $L_{2}$, $\epsilon=3.0$ $\epsilon=4/255$ ($L_{\infty}$, stage 1), $\epsilon=3.0$ ($L_{2}$, stage 2)
Table 5: Adversarial training parameters for $\mathcal{L}_{\text{BCE}}$ (stage 2 only).
CIFAR-10/100 ImageNet (ResNet/WRN) ImageNet (ConvNeXt-L-CvSt) Max PGD steps ($T$) 40/45/50 15/30/65 110 PGD step size 0.1 2.0 3.0 $L_{2}$ perturbation bound None (unconstrained) None (unconstrained) None (unconstrained) OOD data source 80M Tiny Images [Torralba et al., 2008] Open Images [Krasin et al., 2016] Open Images [Krasin et al., 2016]
B.1.2 Data augmentation
As described in Section˜3.6, we implement separate data augmentation pipelines for the discriminative and generative components of our objective function. Table˜6 summarizes these augmentation strategies for stage 2 training. For CIFAR-10/100 and ImageNet ResNet/WRN, the pretrained models we initialize from were trained with the same augmentations as stage 2 $\mathcal{L}_{\text{AT-CE}}$ (from Augustin et al. [2020] for CIFAR and Salman et al. [2020] for ImageNet). For ImageNet ConvNeXt-L-CvSt, the pretrained model from Singh et al. [2023] was trained with heavy augmentations (RandAugment + MixUp + CutMix + Random Erasing), which differ from the simpler stage 2 $\mathcal{L}_{\text{AT-CE}}$ strategy shown in Table˜6. The effects of CIFAR-10 augmentations are illustrated in Figure˜3.
Figure˜4 illustrates CIFAR-10 training curves from stage 2 joint training with various augmentation strategies applied to $\mathcal{L}_{\text{BCE}}$ (while consistently using AutoAugment with Cutout for $\mathcal{L}_{\text{AT-CE}}$). Interestingly, the choice of augmentation for the generative component influences discriminative performance as well, as evidenced by the decline in robust test accuracy when using no augmentation. The best FID performance is achieved by no augmentation and random cropping with padding, which minimally distort the underlying data distribution $p_{\text{data}}$. Overall we find random cropping with padding provides the optimal balance between discriminative and generative performances.
Table 6: Data augmentation strategies for discriminative and generative components.
Dataset Component Augmentation Strategy CIFAR-10/100 $\mathcal{L}_{\text{AT-CE}}$ AutoAugment + Cutout + RandomHorizontalFlip() $\mathcal{L}_{\text{BCE}}$ RandomCrop(32, padding=4) + RandomHorizontalFlip() ImageNet $\mathcal{L}_{\text{AT-CE}}$ RandomResizedCrop(256) + RandomHorizontalFlip() $\mathcal{L}_{\text{BCE}}$ Resize(256) + CenterCrop(256) + RandomHorizontalFlip()
Figure 3: Samples produced by different augmentations on CIFAR-10.
Figure 4: CIFAR-10 training curves under different data augmentations during stage 2 joint training.
B.2 Computational efficiency
Training overhead relative to standard AT. We estimate computational cost as follows: a forward pass costs 1 unit, an input-gradient backward (for PGD) costs 1 unit, and a full backward (for parameter updates) costs 2 units. This gives 2 units per PGD step and 3 units per training update. Under this model, Stage 1 (standard AT) costs $(2K+3)B$ units per iteration: $2KB$ for the $K$-step PGD attack plus $3B$ for the training update. Stage 2 adds $T$ generative PGD steps ($2TB$ units) and processes three sample types per iteration (adversarial in-distribution, adversarial OOD, and clean in-distribution), tripling the training update cost to $9B$ units, for a total of $(2K+2T+9)B$ units per iteration. With $E_{1}$ and $E_{2}$ epochs for each stage, the total overhead relative to standard AT is:
$\text{Overhead}=1+\frac{E_{2}}{E_{1}}\cdot\frac{2K+2T+9}{2K+3}$ (29)
[…]
Table 7: Computational cost and performance metrics of DAT models. Training overhead computed relative to standard AT using Equation˜29. Stage 2 training times are effective durations excluding evaluation. Baseline training times converted from V100-days; throughput measured on a single MI300 accelerator.
[…]
B.3 Training dynamics
We present the stage 2 training curves to illustrate the training dynamics.
(a) CIFAR-10 ($T=40$)
(b) CIFAR-100 ($T=45$)
(c) ImageNet $\times256$ (ResNet-50, $T=15$)
(d) ImageNet $\times256$ (ConvNeXt-L, $T=110$)
Figure 5: Training curves from Stage 2 joint training demonstrating substantial FID score improvements while preserving Stage 1 robust test accuracy (evaluated via PGD attacks; FID measured using 10K generated samples).
B.4 Evaluation methodology
We evaluate generative performance using Fréchet Inception Distance (FID) and Inception Score (IS). FID is computed between 50K class-balanced generated samples and the full training set, while IS is computed on the same set of 50K generated samples.
Conditional generation. We generate an equal number of samples for each class. To generate samples for a given class $y$, we first sample an OOD data point $x$ from the corresponding OOD data source, and then perform $T$ steps of PGD attack according to:
[…]
where $T$ is the number of PGD steps and $\eta$ is the corresponding step size (see Table˜8).
Unconditional generation. For unconditional generation, we directly sample from the marginal distribution using PGD according to Equation˜8:
[…]
The FID results for both conditional and unconditional generation across all datasets are presented in Table˜9. We find conditional generation consistently outperforms unconditional generation across all the datasets.
Table 8: Sample generation parameters for FID and IS evaluation. The number of PGD steps for each model and dataset combination is determined through grid search.
Model Dataset PGD steps ($T$) Step size OOD data source DAT CIFAR-10 ($T=40$) 33 0.2 80M Tiny Images CIFAR-10 ($T=50$) 35 0.2 80M Tiny Images CIFAR-100 ($T=45$) 32 0.2 80M Tiny Images CIFAR-100 ($T=50$) 33 0.2 80M Tiny Images ImageNet (ResNet-50, $T=15$) 14 8.0 Open Images ImageNet (ResNet-50, $T=30$) 14 8.0 Open Images ImageNet (WRN-50-4, $T=30$) 17 8.0 Open Images ImageNet (WRN-50-4, $T=65$) 19 8.0 Open Images ImageNet (ConvNeXt-L-CvSt, $T=110$) 36 8.0 Open Images RATIO CIFAR-10 31 0.2 80M Tiny Images CIFAR-100 14 0.2 80M Tiny Images Standard AT CIFAR-10 22 0.2 80M Tiny Images CIFAR-100 15 0.2 80M Tiny Images ImageNet (ResNet-50) 13 8.0 Open Images ImageNet (WRN-50-4) 11 8.0 Open Images ImageNet (ConvNeXt-L-CvSt) 0 8.0 Open Images
Table 9: FIDs of conditional and unconditional generation of our approach.
CIFAR-10 CIFAR-100 ImageNet $\times224$ (ResNet50, $T=15$) Conditional generation $9.07$ $10.70$ 6.64 Unconditional generation $20.57$ $13.56$ 18.67
Appendix C Results & Analyses
C.1 Qualitative samples
(a) Seed images used for producing the generated samples.
(b) Uncurated conditional samples of DAT ($T=50$).
(c) Uncurated conditional samples of RATIO.
Figure 6: CIFAR-10 class-conditional generation results. Note that some samples from the RATIO baseline show potential artifacts (e.g., saturated or unnatural colors) possibly linked to the aggressive AutoAugment policy used for model training.
(a) Seed images used for producing the generated samples.
(b) Uncurated conditional samples of DAT ($T=50$).
(c) Uncurated conditional samples of RATIO.
Figure 7: CIFAR-100 conditional generation results.
(a) Seed images used for producing the generated samples.
(b) Uncurated conditional samples of DAT (WRN-50-4 $T=65$).
(c) Uncurated conditional samples of standard AT (WRN-50-4).
Figure 8: ImageNet class-conditional generation results for the first 10 classes: tench, goldfish, great white shark, tiger shark, hammerhead, electric ray, stingray, cock, hen, ostrich (images are in $\times256$ resolution).
Figure 9: Selected ImageNet conditional generation results for class 88 (macaw), 107 (jellyfish), 130 (flamingo), 145 (king penguin), 248 (husky), 258 (Samoyed), 291 (lion), 511 (convertible), and 980 (volcano). Results are generated with DAT ConvNeXt-L-CvSt at $\times256$
C.2 Visual counterfactuals
Figure 10: CIFAR-10 counterfactual examples with perturbation limits of $0.5,1.0,1.5,2.0,2.5,3.0$. These figures display counterfactuals and corresponding classifier confidences for both the correct class (top row) and a target wrong class (bottom row). As the perturbation budget increases from left to right, the generated counterfactuals progressively resemble samples from the target class distribution while the target class confidence correspondingly increases, demonstrating that our model effectively captures the distributions of different classes and can generate meaningful class-to-class transformations.
Figure 11: ImageNet counterfactual examples with perturbations limits of $10.,20.,30.,40.,50$.
C.3 OOD detection
We evaluate both standard out-of-distribution (OOD) detection performance and worst-case OOD detection under adversarial perturbations using models trained with $\times224$ generation. For standard OOD detection, we measure the AUROC scores between in-distribution test samples and unmodified OOD samples. For worst-case detection, we evaluate against adversarially perturbed OOD samples specifically optimized to maximize the OOD detection function output. Results are computed using all the in-distribution test samples and 1024 out-distribution samples. For generating adversarial OOD samples, we use $L_{2}$-based perturbation limit of 1.0 for CIFAR-10/100 and 3.0 for ImageNet.
Energy-based detection. We use an energy-based function $s_{\theta}(x)=-E_{\theta}(x)$, which is proportional to $\log p_{\theta}(x)$ up to an additive constant. To find adversarial OOD inputs for this function, we employ a PGD attack to maximize the negative energy:
$x_{adv}=\operatorname*{arg\,max}_{x^{\prime}\in B(x,\epsilon_{o})}-E_{\theta}(x^{\prime})$ (31)
where $x$ is a clean OOD input and $B(x,\epsilon_{o})$ represents an $L_{2}$-ball of radius $\epsilon_{o}$ centered at $x$.
Maximum confidence detection. We employ a maximum confidence function $s_{\theta}(x)=\max_{y}p_{\theta}(y|x)$ that uses the confidence in the most likely class (also used by RATIO [Augustin et al., 2020]). For this detection function, following RATIO [Augustin et al., 2020], we compute adversarial OOD inputs by maximizing the cross-entropy loss against a uniform distribution:
$x_{adv}=\operatorname*{arg\,max}_{x^{\prime}\in B(x,\epsilon_{o})}\mathcal{L}_{\text{CE}}(\theta;x^{\prime},\mathbf{1}/K)$ (32)
where $\mathbf{1}/K$ represents a uniform distribution over all $K$ classes. Maximizing this loss encourages the model to produce a non-uniform (confident) prediction, thereby maximizing the detection function.
Table˜10 presents a comparison of the above two OOD detection functions. The results reveal complementary strengths: the energy-based function ($E_{\theta}(x)$) achieves near-perfect AUROC scores on uniform noise detection, while the maximum confidence function ($\max_{y}p_{\theta}(y|x)$) demonstrates superior performance on natural image OOD datasets. Based on these findings, we adopt the maximum confidence score for subsequent comparisons with other methods.
Table˜11 presents comparative results across different baselines. Notably, our DAT model achieves OOD detection performance comparable to standard AT on natural image datasets (CIFAR-100, SVHN), despite incorporating an additional OOD dataset during training. This observation suggests that our generative training component primarily enhances generation quality rather than improving OOD detection capabilities beyond those provided by standard adversarial training.
Compared to RATIO, our model exhibits lower OOD detection performance across most datasets. To investigate whether this gap stems from our use of milder augmentations for the generative component, we trained an ablation model that applies RATIO’s aggressive augmentation strategy to both loss terms. The results show that this variant performs similarly to our standard DAT model and still underperforms RATIO. This finding indicates that the performance gap is not primarily caused by the augmentation strategy but rather by the fundamental differences in the training objectives: RATIO’s loss is explicitly optimizes for OOD detection performance, while our generative loss prioritizes learning an accurate energy function for generation.
A potential avenue for addressing this limitation involves developing a hybrid objective that combines our generative loss with RATIO’s explicit OOD detection term. This approach is theoretically motivated by the complementary nature of these objectives: our generative component learns to model the energy landscape of in-distribution data while naturally assigning low probability to out-of-distribution regions, which aligns conceptually with RATIO’s strategy of enforcing low confidence predictions in neighborhoods around OOD samples. Such a hybrid formulation could potentially preserve the generative modeling benefits of our approach while recovering the superior OOD detection performance of RATIO. Future work could explore this direction by investigating appropriate weighting strategies between the generative and OOD detection terms to achieve optimal performance across both objectives.
Table 10: Comparison of OOD detection functions on CIFAR-10.
CIFAR-100 SVHN Uniform noise Method Clean Adversarial Clean Adversarial Clean Adversarial DAT ($\max_{y}p_{\theta}(y|x)$) 0.8709 $0.6480$ 0.9609 0.8334 $0.8922$ $0.8257$ DAT ($E_{\theta}(x)$) $0.8484$ 0.6647 $0.8011$ $0.6046$ 0.9995 0.9983
Table 11: OOD detection performance (AUROC) with CIFAR-10 as ID dataset (JEM results are from Augustin et al. [2020]). All methods use the maximum confidence detection function $s_{\theta}(x)=\max_{y}p_{\theta}(y|x)$.
CIFAR-100 SVHN Uniform noise Method Clean Adversarial Clean Adversarial Clean Adversarial JEM $0.8760$ $0.1920$ $0.8930$ $0.0730$ $0.1180$ $0.0250$ Standard AT $0.8759$ $0.6364$ $0.9625$ $0.8306$ $0.8501$ $0.7902$ DAT ($T=40$, uniform aug) $0.8751$ $0.6261$ $0.9642$ $0.8303$ $0.9546$ $0.9254$ DAT ($T=40$) $0.8709$ $0.6480$ $0.9609$ $0.8334$ $0.8922$ $0.8257$ RATIO 0.9157 0.7516 0.9843 0.9130 0.9999 0.9999
Table 12: OOD detection performance (AUROC) with CIFAR-100 as ID dataset.
CIFAR-10 SVHN Uniform noise Method Clean Adversarial Clean Adversarial Clean Adversarial Standard AT 0.7430 0.4093 0.8700 0.4863 0.7858 0.5048 RATIO 0.7320 0.3795 0.8439 0.4356 0.7769 0.5881 DAT ($T=45$) 0.7027 0.5145 0.8271 0.5823 0.4024 0.2283
Table 13: OOD detection performance (AUROC) with ImageNet as ID dataset.
CIFAR-10 SVHN Uniform noise Method Clean Adversarial Clean Adversarial Clean Adversarial Standard AT (ResNet-50) 0.7235 0.5304 0.9239 0.8089 0.8678 0.8377 DAT (ResNet-50, $T=15$) 0.6599 0.4870 0.8813 0.7754 0.6899 0.6268
C.4 Calibration
We assess model calibration using reliability diagrams on models trained with $\times224$ generation. Figures 12, 13, and 14 show calibration diagrams for CIFAR-10, CIFAR-100, and ImageNet respectively.
(a) Standard AT
(b) RATIO
(c) DAT
Figure 12: Calibration diagrams on CIFAR-10 (without temperature scaling).
(a) Standard AT
(b) RATIO
(c) DAT
Figure 13: Calibration diagrams on CIFAR-100 (without temperature scaling).
(a) Standard AT (ResNet-50)
(b) DAT (ResNet-50, $T=15$)
(c) DAT (ResNet-50, $T=30$)
(d) Standard AT (WRN-50-4)
(e) DAT (WRN-50-4, $T=30$)
(f) DAT (WRN-50-4, $T=65$)
Figure 14: Calibration diagrams on ImageNet (without temperature scaling).
C.5 Corruption robustness
To evaluate whether the joint training objective maintains robustness under distribution shift, we assess our CIFAR-10 models on CIFAR-10-C [Hendrycks and Dietterich, 2019], a benchmark testing robustness to 15 common corruption types across 5 severity levels.
As shown in Table˜14, standard AT achieves a mean corruption error (mCE) of 19.63%, while DAT with $T=40$ achieves mCE of 19.84%, and DAT with $T=50$ achieves mCE of 21.84%. While there is a trade-off between generative quality and corruption robustness as $T$ increases, DAT overall maintains strong corruption robustness comparable to standard AT across all corruption types.
Table 14: Robustness evaluation on CIFAR-10. We report clean accuracy, robust accuracy, and corruption robustness on CIFAR-10-C. For corruptions, we report error rate (%) averaged across 5 severity levels. mCE denotes mean corruption error across all 15 corruption types.
Metric Standard AT DAT ($T=40$) DAT ($T=50$)
Noise corruptions (error %)
   Gaussian noise 21.30 20.52 20.63
   Shot noise 17.14 16.36 16.86
   Impulse noise 23.80 23.54 24.72
Blur corruptions (error %)
   Defocus blur 11.08 11.69 13.13
   Glass blur 15.38 14.74 17.06
   Motion blur 13.65 14.27 15.84
   Zoom blur 11.72 12.39 13.89
Weather corruptions (error %)
   Snow 12.19 12.19 13.62
   Frost 12.44 12.13 13.60
   Fog 24.75 26.67 29.45
   Brightness 8.26 8.60 9.61
Digital corruptions (error %)
   Contrast 33.31 32.89 36.24
   Elastic transform 12.34 13.06 14.80
   Pixelate 9.51 9.65 11.08
   JPEG compression 9.56 9.73 11.14
Mean corruption error (mCE) $\downarrow$ 19.63 19.84 21.84
Clean Acc (%) $\uparrow$ 92.43 91.86 90.72
Robust Acc (%) $\uparrow$ 75.73 75.66 74.65
FID $\downarrow$ 28.41 9.07 7.19
C.6 L-infinity generalization
All our models in Tables˜1 and 2 are trained with $L_{2}$-based adversarial attacks for both the discriminative and generative components. To demonstrate that our approach generalizes to $L_{\infty}$ adversarial training, we train on ImageNet $\times256$ with ConvNeXt-L using $L_{\infty}$ attacks: for discriminative training, we use 2-step PGD with $\epsilon=4/255$ and step size 2/255; for generative training, we use step size 3/255 with $T=110$ maximum steps. All other hyperparameters (optimizer, learning rate, weight decay, EMA, batch size) match our $L_{2}$ configuration exactly (see Table˜3). FID is evaluated on samples generated using $L_{\infty}$-based PGD with 36 steps and step size 0.03 (approx. 8/255).
Table˜15 presents the results. Compared to Standard AT, our $L_{\infty}$-trained DAT model trades modest reductions in clean accuracy (76.58% vs 78.25%) and robust accuracy (57.94% vs 59.40%) for significantly superior generation quality (FID 4.11 vs 44.46, IS 320.7 vs 27.32). This demonstrates that DAT successfully achieves joint discriminative-generative modeling under $L_{\infty}$ training, maintaining competitive robustness while enabling high-quality generation.
[…]
Table 15: $L_{\infty}$ and $L_{2}$ results on ImageNet $\times256$ with ConvNeXt-L. Standard AT results are based on $L_{\infty}$-trained ($\epsilon=4/255$) checkpoint from Singh et al. [2023]. Both DAT models initialize from the same $L_{\infty}$-trained Stage 1 checkpoint [Singh et al., 2023] (no $L_{2}$-trained checkpoint is available) and differ only in Stage 2 training: one uses $L_{\infty}$ perturbations, the other uses $L_{2}$ perturbations. All models are evaluated under both $L_{\infty}$ ($\epsilon=4/255$) and $L_{2}$ ($\epsilon=3.0$) adversarial attacks using AutoAttack [Croce and Hein, 2020].
Method Training Norm Clean Acc% $\uparrow$ Robust Acc ($L_{\infty}$ 4/255) $\uparrow$ Robust Acc ($L_{2}$ 3.0) $\uparrow$ FID $\downarrow$ IS $\uparrow$
Standard AT $L_{\infty}$ 78.25 59.40 33.38 44.46 27.32
DAT ($T=110$) $L_{\infty}$ 76.58 57.94 33.40 4.11 320.7
DAT ($T=110$) $L_{2}$ 75.73 51.90 56.40 3.29 310.2
C.7 Ablation studies
In this section, we dissect the contributions of individual model components, analyze the data efficiency of our generative learning mechanism, and investigate the inherent trade-off between discriminative and generative performance.
C.7.1 Component ablation
To analyze the individual contribution of our primary contributions, we conduct an ablation study with the following variants on CIFAR-10:
  • Standard AT: A baseline adversarially trained model without a generative component.
  • DAT with uniform augmentation: Our DAT approach that applies the same aggressive augmentation for both the discriminative and generative objectives.
  • DAT with decoupled augmentation: Our DAT approach that applies aggressive augmentation to the discriminative term and mild augmentation for the generative term.
The results in Table˜16 demonstrate the impacts of the AT-based generative loss and decoupled augmentation. Introducing the generative loss component to the standard AT baseline significantly reduces the FID from 33.04 to 15.35, while robust accuracy remains comparable. The subsequent application of a decoupled augmentation strategy yields a further reduction in FID to 9.07.
Since both our approach and RATIO extend a standard AT baseline with an objective function that leverages out-of-distribution (OOD) data, it is instructive to compare their relative efficacy in enhancing generative fidelity. The RATIO objective, which is formulated for robust OOD detection, reduces the FID from 33.04 to 21.96. In contrast, our generative objective provides a much larger improvement, lowering the FID to 15.35. This comparison confirms that for the goal of enhancing sample quality, a dedicated generative loss is more effective than an auxiliary loss designed for OOD detection.
Table 16: Effect of generative loss and augmentation on CIFAR-10.
Method Acc% $\uparrow$ Robust Acc% $\uparrow$ FID $\downarrow$
Standard AT 92.34 75.73 33.04
DAT (uniform aug) 92.68 75.93 15.35
DAT (decoupled aug) 91.86 75.66 9.07
RATIO [Augustin et al., 2020] 92.23 76.25 21.96
C.7.2 OOD data efficiency
The out-of-distribution (OOD) dataset is a critical component of our training framework, as it provides the initialization samples for computing the negative samples in the generative loss term. The influence of this dataset can be understood through the EBM learning mechanism: a more diverse OOD dataset provides better coverage of the input space, allowing the PGD attack (acting as an MCMC sampler) to discover a broader range of spurious modes in the current energy landscape. These discovered modes are then eliminated as the main objective function is optimized. Given this crucial role, the diversity and scale of the OOD dataset are expected to influence model performance.
To examine the impact of OOD dataset size, we conducted an ablation study on ImageNet using our DAT ResNet-50 ($T=15$) model with varying OOD dataset sizes: 1K, 10K, 100K, and the full 300K samples. As shown in Table 17, the FID score improves modestly from 6.96 with 1K samples to 6.64 with 300K samples. Classification accuracy remains stable across all dataset sizes, with similar robustness levels, indicating that the OOD dataset size primarily affects generation quality rather than discriminative performance.
These results demonstrate notable data efficiency, with only modest improvements when scaling from 1K to 300K OOD samples. A contributing factor to this efficiency is data augmentation: we employ RandomResizedCrop with scale=(0.08, 1.0) and aspect ratio=(0.75, 1.33), which can crop as little as 8% of the original image with varying aspect ratios, potentially amplifying the effective diversity of each sample. To investigate the contribution of augmentation, we include a baseline using 1K OOD samples without data augmentation. While augmentation provides clear benefits—improving FID from 8.00 to 6.96—even without augmentation, our approach substantially outperforms standard AT in generation quality (FID 8.00 vs. 15.97). This indicates the data efficiency can be largely attributed to the AT-based EBM learning mechanism itself, where the PGD attack can effectively explore the energy landscape even from limited initialization points.
Table 17: Impact of OOD dataset size on ImageNet performance for DAT ResNet-50 ($T=15$) with $\times224$ generation.
Method Acc% $\uparrow$ Robust Acc% $\uparrow$ FID $\downarrow$ IS $\uparrow$
Standard AT 62.83 34.44 15.97 274.90
DAT 1K w/o aug 57.50 33.80 8.00 320.64
DAT 1K 57.56 34.22 6.94 324.23
DAT 10K 57.82 34.70 6.84 320.78
DAT 100K 58.19 34.88 6.70 322.10
DAT 300K 57.88 34.84 6.64 339.55
C.7.3 Discriminative-generative trade-off
Our empirical analyses across multiple dimensions—varying PGD steps $T$ (Section˜4.3), adjusting loss weights (see below), and modifying augmentation strategies (Section˜B.1.2)—reveal an inherent trade-off between generative and discriminative performance. The augmentation ablation provides insight into a potential underlying mechanism: when all other settings remain fixed and only the augmentation strategy for the generative component varies, using no augmentation achieves better FID but worse robust accuracy, while random cropping maintains similar FID while improving robustness. This sensitivity to the generative pipeline’s data distribution suggests that model representations adapt to the data used for generative modeling. The substantially lower FID of DAT compared to standard AT is consistent with the learned representations exhibiting stronger alignment with the original data distribution $p_{\text{data}}$ than with the augmented distribution used for discriminative training. These observations suggest a tension between two objectives: the generative objective encourages representations aligned with $p_{\text{data}}$ for high-fidelity generation, while the discriminative objective benefits from alignment with the augmented distribution for robustness. While this fundamental tension cannot be fully resolved within the current framework, the trade-off can be tuned through mechanisms such as the number of PGD iterations $T$ (Equation˜8) and loss weighting to suit different application requirements.
Control via loss weighting. To explicitly investigate the controllability of this trade-off, we perform experiments on CIFAR-10 with three distinct weighting configurations for the composite objective $\mathcal{L}(\theta)=\lambda_{1}\mathcal{L}_{\text{AT-CE}}(\theta)+\lambda_{2}\mathcal{L}_{\text{BCE}}(\theta)$: (1) standard loss with $\lambda_{1}=\lambda_{2}=1.0$; (2) emphasize generative with $\lambda_{1}=0.6,\lambda_{2}=1.4$; and (3) emphasize classification with $\lambda_{1}=1.4,\lambda_{2}=0.6$. The results in Table 18 confirm that the balance between generative and discriminative performance can be tuned by adjusting the loss term weights. Emphasizing the generative component improves FID at the cost of slightly reduced classification performance, while emphasizing classification achieves the opposite effect. However, we note that our standard, unweighted loss corresponds to the natural factorization of the joint log-likelihood in the original JEM formulation: $\log p_{\theta}(x,y)=\log p_{\theta}(y|x)+\log p_{\theta}(x)$. This suggests that equal weighting is a principled default that performs well without requiring additional hyperparameter tuning.
Table 18: Trading off generative and discriminative performance by weighting loss terms.
Method Acc% $\uparrow$ Robust Acc% $\uparrow$ FID $\downarrow$
Standard loss 91.88 75.73 9.09
Emphasize generative modeling 91.16 75.11 8.77
Emphasize classification 92.52 75.97 10.02
C.8 Reproducibility and stability
Table 19: Mean and standard deviation of DAT performance across datasets, computed over five independent runs with different random seeds. We observed zero divergences across all runs.
Dataset Acc% $\uparrow$ Robust Acc% $\uparrow$ IS $\uparrow$ FID $\downarrow$ CIFAR-10 (WRN-34-10, $T=40$) 91.92 $\pm$ 0.09 75.75 $\pm$ 0.07 9.92 $\pm$ 0.05 9.12 $\pm$ 0.05 CIFAR-100 (WRN-34-10, $T=45$, LR=0.009) 65.76 $\pm$ 0.75 45.94 $\pm$ 0.48 10.99 $\pm$ 0.29 10.73 $\pm$ 0.25 ImageNet $\times256$ (ResNet-50, $T=15$) 61.31 $\pm$ 0.16 39.96 $\pm$ 0.41 322.65 $\pm$ 2.28 6.87 $\pm$ 0.05